Hi i have a few questions.
1. which is the larger number 2.0x10^2 or 2.0x10^-3?
2.Convert 27% to a FRACTION
3. If PV=gRT/M evaluate M
4. If d=1.5g/ml and m=3.0g find v in d=m/v
5.If P1/T1=P2/T2 evaluate T2
6. The number 32 is two thirds of what number?
7. Convert 151K to Celcius using formula K=273+celcius

Answer by funmath(2873)

1. 2x10^2=200 and 2x10^-3=.002, 200 is larger than .002, so 2x10^2 is larger than 2X10^-3
----------------------------------------------------------------------------------
2. 27%=27/100
----------------------------------------------------------------------------------
3. PV=gRT/M
M(PV)=M(gRT/M)
MPV=gRT
MPV/PV=gRT/PV
M=gRT/PV
----------------------------------------------------------------------------------
4. d=m/v
v(d)=v(m/v)
vd=m
vd/d=m/d
v=m/d
m=3.0g and d=1.5g/ml substitute these values into the formula
v=3.0g/1.5g/ml
v=2ml
--------------------------------------------------------------------------------
5. P1/T1=P2/T2 Cross multiply.
P1T2=P2T1
P1T2/P1=P2T1/P1
T2=P2T1/P1
--------------------------------------------------------------------------------
6. 32=(2/3)x
(3/2)(32)=(3/2)(2/3)x
96/2=6/6x
48=x
--------------------------------------------------------------------------------
7. 151=273+C
-273+151=-273+273+C
-122=C
C=-122 degrees celsius